# Metric Graph Kernels via the Tropical Torelli Map

### v1

@@ 3.3 Generic Length Functions
Definition 3.
Let $G=(V,E)$ be a connected graph. A length function $\ell:E:\to\mathbb{R}_{+}$ is generic if $G$ has a unique minimal spanning tree $T$ and the lengths of edges are distinct.
If $G$ is equipped with a generic length function, it admits a canonical orientation. We prove our main theorem on uniqueness and refinement-invariance of the resulting matrix under the tropical Torelli map for weighted graphs in Appendix A.2.
Theorem 1.
@@ 4 Tropical Metric Graph Kernels
A common approach to constructing a kernel function on a geometric or structured space is to start with a distance function that captures meaningful dissimilarities between elements. Given a distance $d(x,y)$ on the space, a typical strategy is to convert it into a similarity measure using a radial basis function (RBF), $k(x,y)=\exp(-\gamma d^{2}(x,y))$, where $\gamma>0$ is a scaling parameter [25, 26, 39]. In this section, we construct our metric graph kernel by first enlarging the embedding space to positive semi-definite (PSD) matrices to enable comparisons between graphs of different genus. Then we construct graph kernels using distances on the space of PSD matrices. We present algorithms to compute these kernels and discuss their computational complexity.
4.1 The Space of Positive Semi-Definite Matrices
@@
The tropical Torelli map embeds a graph of genus $g$ as a $g\times g$ symmetric positive definite (SPD) matrix. To compare graphs with different genus, we map the matrices to the same dimension. This is done by modifying the SPD matrices in the following way: Given an upper bound $g_{0}$, we either (i) pad a matrix $Q$ by zero if $\dim(Q)<g_{0}$, or (ii) extract a submatrix by randomly choosing $g_{0}$ rows and columns if $\dim(Q)>g_{0}$. In this way, we obtain a map from graph space to the space of positive semi-definite (PSD) matrices $\mathrm{PSD}(g_{0})$.
@@
Unlike the space of SPD matrices, which forms a smooth manifold that admits various Riemannian metrics, the space of PSD matrices lacks a manifold structure due to the presence of boundary points corresponding to rank-deficient matrices. It is known that $\mathrm{PSD}(g_{0})$ is a stratified space which contains $\mathrm{SPD}(g_{0})$ as its maximal strata [31, 35].
The Tropical Torelli–Euclidean Kernel. We begin by introducing a graph kernel based on the Euclidean distance which is a straightforward approach and provides a simple starting point. Let $G_{1}$ and $G_{2}$ be two weighted graphs, we define the tropical Torelli–Euclidean (TTE) kernel as follows:
@@
where $\|\cdot\|_{F}$ is the Frobenius norm of matrices.
Although $k_{\mathrm{TTE}}$ does not take into account the inherent geometry of $\mathrm{PSD}(g_{0})$, it has two key advantages: first, the computation is efficient, and second, the kernel induced by the Euclidean distance is known to be positive definite [12]. See Appendix B.1 for a full discussion.
4.2 The Tropical Torelli–Wasserstein Kernel
We now turn to proposing a geometric distance on the space of PSD matrices. In the original definition of the tropical Torelli map, a metric graph is sent to a $g$-dimensional flat torus—the quotient space of $\mathbb{R}^{g}$ by the lattice spanned by the column vectors of $\sqrt{Q}$. Since the flat tori are invariant under orthogonal actions, we consider the following distance on $\mathrm{PSD}(g_{0})$:
$W(Q(G_{1}),Q(G_{2}))$ $=\min_{U_{1},U_{2}\in O(g_{0})}\big{\|}U_{1}\sqrt{Q(G_{1})}-U_{2}\sqrt{Q(G_{2})}\big{\|}_{F}$ (2)
$=\min_{U\in O(g_{0})}\big{\|}U\sqrt{Q(G_{1})}-\sqrt{Q(G_{2})}\big{\|}_{F},$
where $O(g_{0})$ is the group of orthogonal matrices. It turns out that (2) is the Bures–Wasserstein distance from information geometry and can be computed from the following closed form [2, 32, 1]:
$W^{2}(Q(G_{1}),Q(G_{2}))=\mathrm{Tr}(Q(G_{1}))+\mathrm{Tr}(Q(G_{2}))-2\mathrm{Tr}\bigg{(}\sqrt{\sqrt{Q(G_{1})}Q(G_{2})\sqrt{Q(G_{1})}}\bigg{)}.$ (3)
On $\mathrm{SPD}(g_{0})$, the Bures–Wasserstein distance corresponds to a Riemannian metric [38, 16, 13]. Though $\mathrm{SPD}(g_{0})$ can be equipped with various different Riemannian metrics, such the log-Euclidean metric and the affine-invariant metric, not all distance functions can be extended to $\mathrm{PSD}(g_{0})$ due to the presence of singularities at the boundary points. We define the tropical Torelli–Wasserstein (TTW) kernel as follows:
$k_{\mathrm{TTW}}(G_{1},G_{2})=\exp(-\gamma W^{2}(Q(G_{1}),Q(G_{2}))).$
As a consequence of Theorem 1, the TTE kernel and TTW kernels are invariant to graph refinement, thus, they are well-defined metric graph kernels. We summarize the computation of the TTE kernel and the TTW kernel in Algorithm 2.
1
2
Input: list of $N$ weighted graph $G$, genus bound $g_{0}$
Output: kernel matrix $K$
3 for $i\leftarrow 1$ to $N$ do
4       compute the tropical Torelli matrix $Q_{i}$ for $G_{i}$
5       if $\mathrm{dim}(Q_{i})<g_{0}$ then
6             $Q_{i}\leftarrow\mathrm{diag}\{Q_{i},0\}$
7
8       else
9             $Q_{i}\leftarrow Q_{i}[\mathrm{rand}(g_{0}),\mathrm{rand}(g_{0})]$
10
11       end if
12
13 end for
[…]
@@ 4.3 Computational Complexity
For a list of $N$ matrices $Q_{i}$ of dimension $g_{0}$, the time complexity of computing the full kernel matrix is $O(N^{2}g_{0}^{2})$ for the TTE kernel and $O(N^{2}g_{0}^{3})$ for the TTW kernel if we use eigenvalue decomposition to compute the square root of a matrix. In practice, we find the computation of the TTE kernel matrix is more efficient for sparse graphs since the matrices $Q_{i}$ can be stored in sparse form which has less memory cost than $O(g_{0}^{2})$.
@@ 5 Experiments
We evaluate our tropical graph kernels with with three experimental studies: (i) Simulations: We simulate graphs of different sparsity and test the practical computational time. (ii) Comparisons: We compare the performance of our kernels with classical label-free graph kernels on standard benchmark datasets. (iii) Real data application: We construct datasets of local urban road networks and demonstrate the applicability of our methods on classification tasks.
@@
Implementation and Computing Infrastructure. We use the open source Python library GraKeL [29], which includes all the graph kernels used in in Section 5.2. Unless otherwise specified, we always use the default hyperparameters provided by GraKeL. All our experiments are performed on a shared server running with 8 CPUs (Intel Xeon Platinum 8358 (Ice Lake) 2.60GHz), 128 GB of RAM, and a walltime limit of 8 hours.
5.1 Verifying Computation Time on Synthetic Datasets
Given node number $n$ and genus $g$, we first generated a line graph with $n$ nodes, and then randomly connected two nodes until the edge number achieves $m=g+n-1$. We set the range of $n$ from $50$ to $140$. For each $n$, we generated $N=10$ synthetic graphs corresponding to each of the following settings of sparsity: (i) Sparse graphs such that $g=c_{s}$ for $c_{s}=5,10,20$; (ii) Semi-sparse graphs such that $g=\lfloor c_{ss}n\rfloor$ for $c_{ss}=1.0,1.5,2.0$; and (iii) Dense graphs such that $g=\lfloor n^{1+c_{d}}\rfloor$ for $c_{d}=0.1,0.2,0.3$. On each synthetic dataset, we repeated the computation of kernel matrices 10 times. Figure 2 shows how the computation time of the TTE and TTW kernels scales with the number of nodes $n$ across datasets with varying sparsity levels. We observe that the practical computation time for the TTE kernel closely follows the growth pattern of the graph genus, whereas the TTW kernel shows faster growth for denser graphs. Moreover, we see that the computation time for both kernels remains nearly constant on sparse datasets regardless of the change of number of nodes, which highlights their potential for applications involving large sparse graphs.
5.2 Comparing to Existing Kernels: Graph Classification on Benchmark Datasets
We now compare the performance of our graph kernels with classical label-free graph kernels on benchmark datasets for graph classification tasks.
Label-Free Graph Kernels. Given that metric graph kernels operate on weighted unlabeled graphs, we focus on classical graph kernels that are likewise independent of node/edge labels and attributes. In particular, we select kernels which can capture global graph structure and are naturally suited to weighted settings. For certain kernels which can be applied to both labeled or unlabeled graphs, we set constant node/edge labels to provide a fair comparison, ensuring that observed performance differences arise from differences in the underlying graph geometry and topology. We compare our kernels with the following five kernels: Edge Histogram (EH) kernel [30]; the Graphlet Sampling (GS) kernel [28]; the ODD-STh (OS) kernel [11]; the Shortest Path (SP) kernel [5]; the $k$-Core Decomposition ($k$CD) Kernel [20]. A detailed description of these kernels and comparison their computational complexity are given in Appendix B.2.
(a) AIDS
(b) FRANKENSTEIN
(c) IMDB-BINARY
(d) MUTAG
(e) NCI109
(f) PROTEIN
(g) REDDIT-BINARY
(h) REDDIT-5K
@@ 5.2 Comparing to Existing Kernels: Graph Classification on Benchmark Datasets
Figure 3: Classification accuracy of 10-fold cross-validation on benchmark datasets (partial).
@@
Benchmark Graph Datasets. We compare the performance of graph kernels on 23 benchmark graph datasets taken from TUDataset [18]. To adapt to the setting of metric graphs, we initialize the node/edge labels and weights of graphs by replacing the original node/edge labels by constants. For unweighted graphs, we add random weights to edges from the uniform distribution $\mathrm{Unif}(0,1)$. To compare computational time, we classify the sparsity of benchmark datasets as follows: Let $\bar{n}$ and $\bar{g}$ be the average number of nodes and genus of a dataset. We identify a dataset as (i) sparse, if $\log(\bar{g}/\bar{n})<0$; (ii) semi-sparse, if $10\leq\log(\bar{g}/\bar{n})<$ 1; (iii) dense, if $\log(\bar{g}/\bar{n})\geq 1$. A complete description of dataset statistics can be found in Table 4 in Appendix C.1.
@@
Test Results. We performed classification using the C-SVM solver in sklearn to test the graph kernels. We used the default hyperparameters in sklearn and evaluated the SVM model through a 10-fold cross-validation. Figure 3 presents the classification accuracy on 8 datasets; a complete list of classification accuracy can be found in Table 5 in Appendix C.2. We see that the TTE and TTW kernels outperform other graph kernels on most benchmark datasets. Since we excluded the label information from the graphs, the classification accuracy of some classical graph kernels are lower than the results reported in previous studies [4, 15].
Table 1: Computation time for benchmark datasets (partial). The EH kernel is excluded.
Name $\log(\bar{g}/\bar{n})$ Sparsity GS OS SP $k$CD TTE TTW
AIDS -2.26 S 9.59 44.30 106.51 231.80 2.53 175.45
BZR -2.29 S 8.57 6.49 3.63 11.81 0.73 9.70
MSRC-9 0.36 SS 111.21 5.99 2.78 12.86 1.85 31.72
MSRC-21 0.45 SS 1004.74 109.10 68.17 257.27 25.97 1497.16
BZR-MD 2.26 D 2328.25 0.97 0.91 63.16 19.21 1410.91
ER-MD 2.31 D 5503.19 1.48 2.78 119.88 31.80 5681.93
@@
We also tested the computation time of feature maps (.fit function in GraKeL) and full kernel matrices (.transform function in GraKeL). We excluded the EH kernel since it is the simplest and fastest. Table 1 presents the total computational time of 6 datasets; complete results can be found in Table 6 and 7 in Appendix C.2. We see that the TTE and TTW kernel are particularly efficient on sparse and semi-sparse graph datasets. For the computation of full kernel matrices, TTE is the most efficient method and gives results on all 23 datasets.

### v2

@@ 3.3 Generic Length Functions
Definition 3.1.
Let $G$ be a connected graph. A length function $\ell:E:\to\mathbb{R}_{+}$ is generic if $G$ has a unique minimal spanning tree $T$ and the lengths of edges are distinct.
If $G$ is equipped with a generic length function, it admits a canonical orientation. We prove our main theorem on uniqueness and refinement-invariance of the resulting matrix under the tropical Torelli map for weighted graphs in Appendix C.1.
Theorem 3.2.
@@
A random length function is generic with probability one: suppose $G$ is a connected graph and for each edge $e_{i}\in E$, we independently assign a random variable $x_{i}$ with uniform distribution on $[0,1]$, then we have $\mathbb{P}[\exists i\neq j,x_{i}=x_{j}]=0$. In real-world data such as road networks, this assumption holds naturally: edge lengths are subject to measurement noise and variability, making exact equality between lengths highly unlikely. As a result, the length function in such settings is generic almost surely.
@@ 4 Tropical Metric Graph Kernels
A common approach to constructing a kernel function on a geometric space is to start with a distance function that captures meaningful dissimilarities between elements. Given a distance $d(x,y)$ on the space, a typical strategy is to convert it into a similarity measure using a radial basis function (RBF), $k(x,y)=\exp(-\gamma d^{2}(x,y))$, where $\gamma>0$ is a scaling parameter (Que and Belkin, 2016; Scholkopf et al., 1997; Zhu et al., 2020). We now construct our metric graph kernel by first enlarging the embedding space to positive semi-definite (PSD) matrices to enable comparisons between graphs of different genus. Then we construct graph kernels using distances on the space of PSD matrices. We present algorithms to compute these kernels and discuss their computational complexity.
4.1 The Space of Positive Semi-Definite Matrices.
@@
The tropical Torelli map sends a weighted graph of genus $g$ to a $g\times g$ symmetric positive definite (SPD) matrix. To compare graphs with different genus, we embed the matrices to a common ambient space. Given a sufficiently large upper bound $g_{0}$, this can be done by simply padding each matrix with zeros, which results in a positive semi-definite (PSD) matrix of fixed size. In this way, graphs are mapped into the same space of PSD matrices $\mathrm{PSD}(g_{0})$.
@@
Unlike the space of SPD matrices, which is a smooth manifold that admits various Riemannian metrics, the space of PSD matrices lacks a manifold structure due to the presence of boundary points corresponding to rank-deficient matrices. It is known that $\mathrm{PSD}(g_{0})$ is a stratified space which contains $\mathrm{SPD}(g_{0})$ as its maximal strata (Thanwerdas and Pennec, 2022; Vandereycken et al., 2013). Many well-studied Riemannian distances on the SPD manifold can not be extended naturally to PSD space. For example, the log-Euclidean and affine-invariant metrics are not well-defined on the PSD space (Bonnabel and Sepulchre, 2010).
On the other hand, for computational efficiency, the ambient space should not be of extremely high dimension. To address this issue, we manually set a target dimension $g_{0}$ and seek to reduce the size of any high-dimensional SPD matrix. In principle, the optimal submatrix should be obtained by minimizing the approximation error in the Frobenius norm. However, the exact solution can only be found by enumerating all permutations of rows and columns, which is computational infeasible. Thus we adopt a simple approach by randomly selecting a principal submatrix of size $g_{0}$. Notice that here the goal is to obtain a small matrix of size $g_{0}$ rather than a large matrix of rank $g_{0}$, which excludes standard low-rank approximation techniques based on the Eckart–Young theorem (Eckart and Young, 1936) or the Nyström method (Drineas et al., 2005). Further discussions and empirical results on stability are provided in Appendix D.
4.2 Two Tropical Kernels for Metric Graphs
We construct our metric graph kernels using the RBF approach and thus now turn to proposing a geometric distance on the space of PSD matrices.
The Bures–Wasserstein Distance. In the original definition of the tropical Torelli map, a metric graph is sent to a $g$-dimensional flat torus which is the quotient space of $\mathbb{R}^{g}$ by the lattice spanned by the column vectors of $\sqrt{Q}$. The flat torus can be represented as $(\mathbb{R}^{g}/\sqrt{Q},\|\cdot\|_{2})$, which allows for isometries under orthogonal transformations. That is, given any orthogonal matrix $U\in O(g_{0})$, the tori $\mathbb{R}^{g}/\sqrt{Q}$ and $\mathbb{R}^{g}/(U\sqrt{Q})$ are equivalent up to isometry. Therefore, we consider the following distance for any $Q_{1},Q_{2}\in\mathrm{PSD}(g_{0})$ modulo the orthogonal transformations:
$d_{W}(Q_{1},Q_{2})=\min_{U_{1},U_{2}\in O(g_{0})}\big\|U_{1}\sqrt{Q_{1}}-U_{2}\sqrt{Q_{2}}\big\|_{F}$ (*)
It turns out that ()(*) is exactly the Bures–Wasserstein distance on $\mathrm{PSD}(g_{0})$ and can be computed from the following closed form (Bhatia et al., 2019; Thanwerdas and Pennec, 2023a; Altschuler et al., 2021):
$(*)=\left(\mathrm{Tr}(Q_{1})+\mathrm{Tr}(Q_{2})-2\mathrm{Tr}\bigg(\sqrt{\sqrt{Q_{1}}Q_{2}\sqrt{Q_{1}}}\bigg)\right)^{\frac{1}{2}}.$
Due to page limits, we defer discussing the motivation from tropical geometry to Appendix A.2. For completeness, we also include a derivation of ()(*) in Appendix C.2.
The Tropical Torelli–Wasserstein Kernel. Based on the tropical Torelli map for weighted graphs and the Bures–Wasserstein distance on the space of PSD matrices, we define the tropical Torelli–Wasserstein (TTW) kernel as
$k_{\mathrm{TTW}}(G_{1},G_{2})=\exp(-\gamma d_{W}^{2}(Q(G_{1}),Q(G_{2}))).$
As a consequence of Theorem 3.2, the TTW kernel is invariant to graph refinement, thus, it is a well-defined metric graph kernel.
To further specify what the TTW kernel captures, we require the notion of 2-isomorphism between weighted graphs (Brannetti et al., 2011).
Definition 4.1.
Two weighted graphs $(G_{1},\ell_{1})$ and $(G_{2},\ell_{2})$ are 2-isomorphic if there exists a bijection $\phi:E(G_{1})\to E(G_{2})$ such that $\ell_{2}(\phi(e))=\ell_{1}(e)$ for all $e\in E(G_{1})$ and $\phi$ induces a bijection between cycles of $G_{1}$ and cycles of $G_{2}$.
Roughly speaking, a 2-isomorphism preserves the combinatorial structure of a graph $G$ at the edge level and induces an isometry on the underlying metric space $|G|$. The following theorem shows that the TTW kernel measures the similarity between two metric graphs in terms of this isometry.
Theorem 4.2.
Let $(G_{1},\ell_{1})$ and $(G_{2},\ell_{2})$ be two weighted graphs without bridges and equipped with generic length functions. If $(G_{1},\ell_{1})$ is 2-isomorphic to $(G_{2},\ell_{2})$, then $k_{\mathrm{TTW}}(G_{1},G_{2})=1$; conversely, if $k_{\mathrm{TTW}}(G_{1},G_{2})=1$, then there exist refinements $G_{1}^{\prime}\geq G_{1}$ and $G_{2}^{\prime}\geq G_{2}$ such that $G_{1}^{\prime}$ is 2-isomorphic to $G_{2}^{\prime}$.
The Tropical Torelli–Euclidean Kernel. Although the TTW kernel enjoys a clear theoretical interpretation via Theorem 4.2 and captures underlying tropical geometry by construction, we make the important remark here that the TTW kernel is not positive definite in the kernel sense, since the 2-Wasserstein distance is not conditionally negative definite on $\mathrm{PSD}(g_{0})$ (Appendix B.2). Therefore, we introduce a simple, alternative kernel obtained by replacing the Bures–Wasserstein distance with the Euclidean distance, which we call the tropical Torelli–Euclidean (TTE) kernel:
@@ 4.2 Two Tropical Kernels for Metric Graphs
Similar to the TTW kernel, the TTE kernel is also a valid metric graph kernel.
Although $k_{\mathrm{TTE}}$ does not take into account the inherent geometry of the space of PSD matrices, it has two key advantages: first, the computation is efficient since it does not require computing the matrix square root, and second, a kernel on $\mathbb{R}^{n}$ induced by the Euclidean distance is known to be positive definite (Haasdonk and Bahlmann, 2004), hence we also have the positive definiteness of the TTE kernel.
Theorem 4.3.
Let $\mathcal{M}_{\leq g}$ be the space of weighted graphs with generic length functions of genus no greater than $g$. Then $k_{\mathrm{TTE}}$ is a positive definite kernel on $\mathcal{M}_{\leq g}$.
@@
0: list of $N$ weighted graph $G$, genus bound $g_{0}$
0: kernel matrix $K$
1:for $i\leftarrow 1$ to NN do
2:  compute the tropical Torelli matrix $Q_{i}$ for $G_{i}$
3:  if $\mathrm{dim}(Q_{i})<g_{0}$ then
4:   $Q_{i}\leftarrow\mathrm{diag}\{Q_{i},0\}$
5:  else
6:   $Q_{i}\leftarrow Q_{i}[\mathrm{rand}(g_{0}),\mathrm{rand}(g_{0})]$
7:   $K\leftarrow\exp(-\gamma\cdot\mathrm{pairwise\_dist}^{2}([Q_{i}]))$
8:  end if
9:end for
@@ 4.3 Computational Complexity
The computation of the TTW kernel and the TTE kernel are given in Algorithm 2. For a list of $N$ matrices $Q_{i}$ of dimension $g_{0}$, the time complexity of computing the full kernel matrix is $O(N^{2}g_{0}^{2})$ for the TTE kernel and $O(N^{2}g_{0}^{3})$ for the TTW kernel if we use eigenvalue decomposition to compute the square root of a matrix. In practice, we find the computation of the TTE kernel matrix is more efficient for sparse graphs since the matrices $Q_{i}$ can be stored in sparse form which has less memory cost than $O(g_{0}^{2})$.
@@ 5 Experiments
We evaluate our tropical graph kernels with three experimental studies: (i) Simulations: We simulate graphs of different sparsity and test empirical computational time and kernel stability. (ii) Comparisons: We compare the performance of our kernels with classical label-free graph kernels on de-labeled benchmark datasets. (iii) Real data application: We construct datasets of local urban road networks and demonstrate the applicability of our methods on classification tasks.
@@
Implementation and Computing Infrastructure. We use the open source Python library GraKeL (Siglidis et al., 2020), which includes all the graph kernels used in in Section 5.2. Unless otherwise specified, we always use the default hyperparameters provided by GraKeL. All our experiments are performed on a shared server running with 8 CPUs (Intel Xeon Platinum 8358 (Ice Lake) 2.60GHz), 128 GB of RAM, and a walltime limit of 8 hours. Code is available upon request.
(a) FRANKENSTEIN
(b) IMDB-BINARY
(c) MUTAG
(d) PROTEIN
(e) REDDIT-BINARY
@@
Figure 2: Classification accuracy of 10-fold cross-validation on benchmark datasets without labels (partial).
5.1 Simulations on Synthetic Datasets
Verifying Computation Runtime. Given node number $n$ and genus $g$, we first generated a line graph with $n$ nodes, and then randomly connected two nodes until the edge number achieves m=g+n1m=g+n-1. We set the range of $n$ from $50$ to $140$. For each $n$, we generated $N=10$ synthetic graphs corresponding to each of the following settings of sparsity: (i) Sparse graphs such that $g=c_{s}$ for $c_{s}=5,10,20$; (ii) Semi-sparse graphs such that $g=\lfloor c_{ss}n\rfloor$ for $c_{ss}=1.0,1.5,2.0$; and (iii) Dense graphs such that $g=\lfloor n^{1+c_{d}}\rfloor$ for $c_{d}=0.1,0.2,0.3$. On each synthetic dataset, we repeated the computation of kernel matrices 10 times. Figure 6 in Appendix D.1 shows how the computation time of the TTE and TTW kernels scales with the number of nodes $n$ across datasets with varying sparsity levels. We observe that the practical computation time for the TTE kernel closely follows the growth pattern of the graph genus, whereas the TTW kernel shows faster growth for denser graphs. Moreover, we see that the computation time for both kernels remains nearly constant on sparse datasets regardless of the change of number of nodes, which highlights their potential for applications involving large sparse graphs.
Verifying Empirical Stability. Using the same strategy for generating synthetic graphs, we evaluate the empirical stability of our kernels with respect to the constructions described in Section 4.
First, we test the stability of random truncation of SPD matrices when reducing a large SPD matrix to a fixed size $g_{0}$. The results are shown in Appendix D.2, from which we see the random truncation method does not induce large variations in the resulting kernel values.
Second, we test the positive definiteness of the TTW and TTE kernels by examining the smallest nonzero eigenvalue of the corresponding kernel matrices. The results are shown in Appendix D.3. In particular, the results are consistent with Theorem 4.3 that the TTE kernel is positive definite, and they show that in practice we do not observe significant instability due to the indefinite nature of the TTW kernel.
Finally, we test the sensitivity of the kernel values with respect to perturbations of edge lengths. We assume an additive noise model on edge length, and we measure the difference of kernel matrices under Frobenius norm. The results are given in Appendix D.4 and we see that our kernels are robust under edge length perturbations.
5.2 Comparison to Existing Kernels
We now compare the performance of our graph kernels to that of existing graph kernels for graph classification tasks. Since ours are the first graph kernels designed for metric graphs, there are no prior baselines available in this setting; thus, we compare against classical label-free graph kernels, on benchmark datasets with labels removed, which represent the most relevant existing alternatives for comparison.
Label-Free Graph Kernels. Given that metric graph kernels operate on weighted unlabeled graphs, we focus on classical graph kernels that are likewise independent of node/edge labels and attributes. In particular, we select kernels that can capture global graph structure and are naturally suited to weighted settings. For certain kernels which can be applied to both labeled or unlabeled graphs, we set constant node/edge labels to provide a fair comparison, ensuring that observed performance differences arise from differences in the underlying graph geometry and topology.
We compare our kernels against the following six label-free graph kernels: the Edge Histogram (EH) kernel (Sugiyama and Borgwardt, 2015), the Graphlet Sampling (GS) kernel (Shervashidze et al., 2009), the ODD-STh (OS) kernel (Da San Martino et al., 2012), the Shortest Path (SP) kernel (Borgwardt and Kriegel, 2005), the Weisfeiler–Lehman (WL) kernel (Shervashidze et al., 2011), and the $k$-Core Decomposition (kkCD) kernel (Nikolentzos et al., 2018). The widely used WL kernel is based on iterative refinement of node labels and therefore is not designed to respect invariances such as edge subdivision that are central to our setting of metric graphs. As a result, WL addresses a different comparison regime than the one we study. Nevertheless, due to its prominence in the literature, we include the WL kernel in our experimental evaluation using a constant initial node labeling. A detailed description of all comparison kernels is provided in Appendix E.1.
@@ 5.2 Comparison to Existing Kernels
Benchmark Graph Datasets without Labels. We compare the performance of graph kernels on 23 benchmark graph datasets taken from the TUDataset (Morris et al., 2020). To adapt to the setting of metric graphs, we initialize the node/edge labels and weights of graphs by replacing the original node/edge labels by constants. For unweighted graphs, we add random weights to edges from the uniform distribution $\mathrm{Unif}(0,1)$. To compare computational time, we classify the sparsity of benchmark datasets as follows: Let $\bar{n}$ and $\bar{g}$ be the average number of nodes and genus of a dataset. We identify a dataset as (i) sparse, if $\log(\bar{g}/\bar{n})<0$; (ii) semi-sparse, if $0\leq\log(\bar{g}/\bar{n})<1$; (iii) dense, if $\log(\bar{g}/\bar{n})\geq 1$. A complete description of dataset statistics can be found in Table 5 in Appendix E.3.
@@
Test Results. We performed classification using the C-SVM solver in sklearn to test the graph kernels. We used the default hyperparameters in sklearn and evaluated the SVM model through a 10-fold cross-validation. Figure 2 presents the classification accuracy on 5 sample datasets; a complete list of classification accuracy can be found in Table 6 in Appendix E.4. We see that the TTE and TTW kernels outperform other graph kernels on most benchmark datasets. Since we excluded the label information from the graphs, the classification accuracy of some classical graph kernels are lower than the results reported in previous studies (Borgwardt et al., 2020; Kriege et al., 2020).
@@
We also tested the computation time of feature maps (.fit function in GraKeL) and full kernel matrices (.transform function in GraKeL). We excluded the EH kernel since it is the simplest and fastest. Complete results of computational time can be found in Table 7(a) and 7(b) in Appendix E.4. We see that the TTE and TTW kernel are particularly efficient on sparse and semi-sparse graph datasets. For the computation of full kernel matrices, TTE is the most efficient method and gives results on all 23 datasets.